# Koby Hayashi

### v1

@@ 4.1.1 Complexity
At a high level the main computational kernels of SymNMF via regularization include ${\bm{\mathbf{{X}}}}{\bm{\mathbf{{H}}}}$, ${\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{X}}}}$, ${\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{W}}}}$, and ${\bm{\mathbf{{H}}}}^{\sf T}{\bm{\mathbf{{H}}}}$, which cost $O(mnk)$ flops for the products with ${\bm{\mathbf{{X}}}}$ and $O(mk^{2})$ for the Gramians. Once this is done the these matrix products are used to perform an update, e.g. via (approximately) solving the NLS problem.
@@
As previously stated, sampling does not generally affect the cost of the (approximate) NLS solve which costs $O(T(m,k))$. The two update rules we use are the HALS and BPP methods. For a discussion of these rules and their properties see [23, 16].
@@
1: data matrix ${\bm{\mathbf{{X}}}}\in\mathbb{R}^{m\times m}$, target rank $k$, threshold $\tau$
2:$\{{\bm{\mathbf{{W}}}},{\bm{\mathbf{{H}}}}\}$ as the factors for an approximate rank-$k$ NMF of ${\bm{\mathbf{{X}}}}$.
@@
10:         ${\bm{\mathbf{{X}}}}_{H}:={\bm{\mathbf{{H}}}}^{\sf T}{\bm{\mathbf{{S}}}}_{H}^{\sf T}{\bm{\mathbf{{S}}}}_{H}{\bm{\mathbf{{X}}}}^{\sf T}+\alpha{\bm{\mathbf{{H}}}}$
@@
12:         ${\bm{\mathbf{{W}}}}:=\text{Update}\Big{(}{\bm{\mathbf{{G_{H}}}}},{\bm{\mathbf{{X}}}}_{H}\Big{)}$
@@
17:         ${\bm{\mathbf{{X}}}}_{W}:={\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{S}}}}_{W}^{\sf T}{\bm{\mathbf{{S}}}}_{W}{\bm{\mathbf{{X}}}}+\alpha{\bm{\mathbf{{W}}}}^{\sf T}$
@@
19:         ${\bm{\mathbf{{H}}}}:=\text{Update}\Big{(}{\bm{\mathbf{{G_{W}}}}},{\bm{\mathbf{{X}}}}_{W}\Big{)}$
@@ 4.2 Practical Considerations for SymNMF via Leverage Score Sampling
This section describes the implementation of Algorithm LvS-SymNMF. We discuss how the leverage scores are computed and the use of a hybrid leverage score sampling scheme which was introduces in [27]. A theoretical analysis of Hybrid Sampling is given later in Section 4.3.2.
@@
Hybrid leverage score sampling samples a subset of rows deterministically and then randomly samples from the remaining rows. We find that Hybrid Sampling is crucial for good performance in our empirical results and we offer a rigorous analysis of its theoretical performance in Section 4.3.2. Hybrid Sampling was proposed and shown to be effective for computing CP decompositions of sparse tensors by Larsen and Kolda [26]. A similar method has been used and theoretically analyzed before for the column subset selection problem [36].
In Hybrid Sampling, a threshold $\tau\in[0,1]$ is used as a hyperparameter. When sampling according to the leverage score distribution, all rows that satisfy $p_{i}\geq\tau$ are deterministically selected. Let the full set of row indices be $\mathscr{I}$, the set that is deterministically sampled be $\mathscr{I}_{D}$ and $s_{D}=|\mathscr{I}_{D}|$, and the rest be $\mathscr{I}_{R}=\mathscr{I}\setminus\mathscr{I}_{D}$ with $s_{R}=|\mathscr{I}_{R}|$, the remaining indices from which random samples are drawn.
Let ${\bm{\mathbf{{Q}}}}_{A}$ be a $m\times k$ matrix with othonormal columns that is being sampled. Assume without loss of generality that the rows of ${\bm{\mathbf{{Q}}}}_{A}$ are sorted in non-increasing order by leverage score. The hybrid sampling matrix takes the form
${\bm{\mathbf{{S}}}}_{H}=\begin{bmatrix}{\bm{\mathbf{{S}}}}_{D}&{\bm{\mathbf{{0}}}}_{s_{D}\times(m-s_{D})}\\ {\bm{\mathbf{{0}}}}_{(s_{R})\times s_{D}}&{\bm{\mathbf{{S}}}}_{R}\end{bmatrix}\in\mathbb{R}^{s_{H}\times m}$ (25)
where ${\bm{\mathbf{{S}}}}_{R}\in\mathbb{R}^{s_{R}\times(m-s_{D})}$ and ${\bm{\mathbf{{S}}}}_{D}\in\mathbb{R}^{s_{D}\times s_{D}}$ is a permutation matrix for the deterministic sampling portion defined as
$({\bm{\mathbf{{S}}}}_{D})_{ji}=\begin{cases}1,\ \text{if row $i$ is the $j$th deterministic sample}\\ 0,\ \text{otherwise}.\end{cases}$ (26)
${\bm{\mathbf{{S}}}}_{D}$ is often included in notation to make the deterministic sampling aspect of equations explicit. The submatrix ${\bm{\mathbf{{S}}}}_{R}$ is a leverage score sampling matrix as defined in Equation 19 but just of the indices in $\mathscr{I}_{R}$. When sampling for ${\bm{\mathbf{{S}}}}_{R}$, rows that were sampled during the deterministic phase are not considered and the leverage scores are renormalized appropriately. The new leverage scores probabilities are $\tilde{p}_{i}=\frac{l_{i}}{k-\theta}$ where $\theta=\sum_{i\in\mathscr{I}_{D}}l_{i}({\bm{\mathbf{{A}}}})$.
@@ 4.3.1 Leverage Score Sampling for NLS Problems
Error bounds and corresponding sampling complexities for sketching the ordinary least squares (OLS) problem have been previously shown in a number of works [9, 38]. Larsen and Kolda gave a bound and proof structure in their work on computing a randomized CP decomposition [26]. Boutsidis and Drineas [3] considered using the randomized Hadamard transform for solving the NLS problem.
We now introduce generalized error bounds and sampling complexities for the nonnegative least squares problem (NLS),
$\min_{{\bm{\mathbf{{x}}}}\geq 0}\|{\bm{\mathbf{{A}}}}{\bm{\mathbf{{x}}}}-{\bm{\mathbf{{b}}}}\|_{2}.$ (27)
We note that our results hold for all constrained least squares (CLS) problems as far as the problem remains convex. The problem in Equation 27 is a convex optimization problem but unlike the case of OLS, does not yield a closed form solution. We are concerned only with the case where ${\bm{\mathbf{{A}}}}$ is full rank and overdetermined. The NLS error bound and sample complexity can be derived based on results by Daniels for the perturbation of Convex Quadratic Programs [7] and leverage score sampling for the OLS problem [9, 33, 27, 45].
We present the following theorem :
Teorema 2.
[…]
Also let $\hat{{\bm{\mathbf{{x}}}}}_{nls}=\operatorname*{arg\,min}_{{\bm{\mathbf{{x}}}}\geq 0}\|{\bm{\mathbf{{S}}}}{\bm{\mathbf{{A}}}}{\bm{\mathbf{{x}}}}-{\bm{\mathbf{{S}}}}{\bm{\mathbf{{b}}}}\|_{2}$ be the sampled NLS problem. Then with probability $1-\delta$, the following holds :
where ${\bm{\mathbf{{r}}}}_{nls}={\bm{\mathbf{{A}}}}{\bm{\mathbf{{x}}}}_{nls}-{\bm{\mathbf{{b}}}}$ and $\sigma_{min}({\bm{\mathbf{{A}}}})$ is the minimum singular value of ${\bm{\mathbf{{A}}}}$  [^1] We note that theorem 2 does not include a leverage score “missestimation factor” ($\beta$ in [45]) which is often included in works concerning leverage score sketching for OLS problems [27, 33, 45]. When inexact leverage scores are used for sampling, the missestimation factor gives a measure of how close the inexact leverage scores are to the true leverage scores. We do not thoroughly discuss missestimation factors because we do not use the concept in this work. However, theorem 2 can be easily generalized to incorporate such a factor..
The following proof uses two Structural Conditions (SC’s) such that if both are true then the error bound in Theorem 2 for the NLS problem holds. Next the conditions, e.g. sampling complexity and probability, are discussed under which these SC’s hold.
Let ${\bm{\mathbf{{A}}}}={\bm{\mathbf{{U}}}}_{A}{\bm{\mathbf{{\Sigma}}}}_{A}{\bm{\mathbf{{V}}}}_{A}^{\sf T}$ be the economy SVD of the coefficient matrix in Equation 27 where ${\bm{\mathbf{{U}}}}_{A}\in\mathbb{R}^{m\times n}$, ${\bm{\mathbf{{\Sigma}}}}_{A}\in\mathbb{R}^{n\times n}$, and ${\bm{\mathbf{{V}}}}_{A}\in\mathbb{R}^{n\times n}$. Let ${\bm{\mathbf{{S}}}}$ be a leverage score sampling matrix for ${\bm{\mathbf{{A}}}}$. The first Structural Condition (SC1) is
$1-\epsilon_{s}\leq\sigma_{i}({\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{S}}}}^{\sf T}{\bm{\mathbf{{S}}}}{\bm{\mathbf{{U}}}}_{A})\leq 1+\epsilon_{s}$ (28)
@@
$\|{\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{r}}}}_{nls}-{\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{S}}}}^{\sf T}{\bm{\mathbf{{S}}}}{\bm{\mathbf{{r}}}}_{nls}\|_{2}\leq\sqrt{\epsilon_{r}}\frac{\|{\bm{\mathbf{{r}}}}_{nls}\|_{2}}{\sqrt{2}}$ (29)
for some $\epsilon_{r}\in(0,1)$. The leverage score sketching matrix ${\bm{\mathbf{{S}}}}$ satisfies SC1 and SC2 with high probability (given sufficiently many samples). For SC1 this is shown in [45] and for SC2 it can be shown by using Theorem 8 on the product ${\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{r}}}}_{nls}$ [33]. We include the associated theorems for these statements in Section H.1.
@@
$\min_{{\bm{\mathbf{{x}}}}\in\mathscr{C}}Q({\bm{\mathbf{{x}}}})=\min_{{\bm{\mathbf{{x}}}}\in\mathscr{C}}\frac{1}{2}{\bm{\mathbf{{x}}}}^{\sf T}{\bm{\mathbf{{K}}}}{\bm{\mathbf{{x}}}}-{\bm{\mathbf{{x}}}}^{\sf T}{\bm{\mathbf{{d}}}}$ (30)
@@
$\min_{{\bm{\mathbf{{x}}}}\geq 0}\|{\bm{\mathbf{{S}}}}{\bm{\mathbf{{A}}}}{\bm{\mathbf{{x}}}}-{\bm{\mathbf{{S}}}}{\bm{\mathbf{{b}}}}\|_{2}$ (31)
@@
$\min_{{\bm{\mathbf{{x}}}}\geq 0}Q({\bm{\mathbf{{x}}}})=\min_{{\bm{\mathbf{{x}}}}\geq 0}\frac{1}{2}{\bm{\mathbf{{x}}}}^{\sf T}{\bm{\mathbf{{A}}}}^{\sf T}{\bm{\mathbf{{A}}}}{\bm{\mathbf{{x}}}}-{\bm{\mathbf{{x}}}}^{\sf T}{\bm{\mathbf{{A}}}}^{\sf T}{\bm{\mathbf{{b}}}},$ (32)
$\min_{{\bm{\mathbf{{x}}}}\geq 0}\hat{Q}({\bm{\mathbf{{x}}}})=\min_{{\bm{\mathbf{{x}}}}\geq 0}\frac{1}{2}{\bm{\mathbf{{x}}}}^{\sf T}{\bm{\mathbf{{A}}}}^{\sf T}{\bm{\mathbf{{S}}}}^{\sf T}{\bm{\mathbf{{S}}}}{\bm{\mathbf{{A}}}}{\bm{\mathbf{{x}}}}-{\bm{\mathbf{{x}}}}^{\sf T}{\bm{\mathbf{{A}}}}^{\sf T}{\bm{\mathbf{{S}}}}^{\sf T}{\bm{\mathbf{{b}}}}.$ (33)
Equation 33 can be interpreted as a perturbed version of Equation 32 with perturbed parameters $\hat{{\bm{\mathbf{{K}}}}}={\bm{\mathbf{{A}}}}^{\sf T}{\bm{\mathbf{{S}}}}^{\sf T}{\bm{\mathbf{{S}}}}{\bm{\mathbf{{A}}}}$ and $\hat{{\bm{\mathbf{{d}}}}}={\bm{\mathbf{{A}}}}^{\sf T}{\bm{\mathbf{{S}}}}^{\sf T}{\bm{\mathbf{{b}}}}$. We make use of the following lemma:
Lemma 3 ([7, Equation 2.4]).
Convex Quadratic Program Inequality: $Q({\bm{\mathbf{{x}}}})$ and $\hat{Q}({\bm{\mathbf{{x}}}})$ from eq. 32 and eq. 33 satisfy
$(\hat{{\bm{\mathbf{{x}}}}}_{nls}-{\bm{\mathbf{{x}}}}_{nls})^{\sf T}\Big{[}\nabla\hat{Q}(\hat{{\bm{\mathbf{{x}}}}}_{nls})-\nabla\hat{Q}({\bm{\mathbf{{x}}}}_{nls})\Big{]}\leq(\hat{{\bm{\mathbf{{x}}}}}_{nls}-{\bm{\mathbf{{x}}}}_{nls})^{\sf T}\Big{[}\nabla Q({\bm{\mathbf{{x}}}}_{nls})-\nabla\hat{Q}({\bm{\mathbf{{x}}}}_{nls})\Big{]}$ (34)
@@
All the tools needed for the proof of Theorem 2 have now been established.
Proof.
Substituting the QP formulations of the original and sampled NLS problems into Equation 34, denoting $\hat{{\bm{\mathbf{{x}}}}}_{nls}-{\bm{\mathbf{{x}}}}_{nls}={\bm{\mathbf{{y}}}}$ and parameterizing in terms of the matrix ${\bm{\mathbf{{U}}}}_{A}$ by writing ${\bm{\mathbf{{Z}}}}={\bm{\mathbf{{\Sigma}}}}_{A}{\bm{\mathbf{{V}}}}^{\sf T}_{A}$ and ${\bm{\mathbf{{Z}}}}{\bm{\mathbf{{y}}}}={\bm{\mathbf{{z}}}}$ the right hand side becomes
@@
${\bm{\mathbf{{y}}}}^{\sf T}{\bm{\mathbf{{Z}}}}^{\sf T}{\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{S}}}}^{\sf T}{\bm{\mathbf{{S}}}}{\bm{\mathbf{{U}}}}_{A}{\bm{\mathbf{{Z}}}}{\bm{\mathbf{{y}}}}={\bm{\mathbf{{z}}}}^{\sf T}({\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{S}}}}^{\sf T}{\bm{\mathbf{{S}}}}{\bm{\mathbf{{U}}}}_{A}){\bm{\mathbf{{z}}}}=\|{\bm{\mathbf{{S}}}}{\bm{\mathbf{{U}}}}_{A}{\bm{\mathbf{{z}}}}\|_{2}^{2}\geq\sigma_{min}({\bm{\mathbf{{S}}}}{\bm{\mathbf{{U}}}}_{A})\|{\bm{\mathbf{{z}}}}\|_{2}^{2}.$
The left hand side of Equation 34 develops to

### v2

@@ 4.1.1 Complexity
At a high level the main computational kernels of SymNMF via regularization include ${\bm{\mathbf{{X}}}}{\bm{\mathbf{{H}}}}$, ${\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{X}}}}$, ${\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{W}}}}$, and ${\bm{\mathbf{{H}}}}^{\sf T}{\bm{\mathbf{{H}}}}$, which cost $O(m^{2}k)$ flops for the products with ${\bm{\mathbf{{X}}}}$ and $O(mk^{2})$ for the Gramians. Once this is done these matrix products are used to perform an update, e.g. via (approximately) solving the NLS problem.
@@
As previously stated, sampling does not generally affect the cost of the Update Rule which costs $O(T(m,k))$. The two update rules we use are the HALS and BPP methods. For a discussion of these rules and their properties see [32, 23].
@@
1:Symmetric data matrix ${\bm{\mathbf{{X}}}}\in\mathbb{R}^{m\times m}$, target rank $k$, threshold $\tau$
@@
10:         ${\bm{\mathbf{{Y}}}}_{H}:={\bm{\mathbf{{H}}}}^{\sf T}{\bm{\mathbf{{S}}}}_{H}^{\sf T}{\bm{\mathbf{{S}}}}_{H}{\bm{\mathbf{{X}}}}^{\sf T}+\alpha{\bm{\mathbf{{H}}}}^{\sf T}$
@@
12:         ${\bm{\mathbf{{W}}}}:=\text{Update}\Big{(}{\bm{\mathbf{{G_{H}}}}},{\bm{\mathbf{{Y}}}}_{H}\Big{)}$triangleright See Appendix E for description of Update()
@@
17:         ${\bm{\mathbf{{Y}}}}_{W}:={\bm{\mathbf{{W}}}}^{\sf T}{\bm{\mathbf{{S}}}}_{W}^{\sf T}{\bm{\mathbf{{S}}}}_{W}{\bm{\mathbf{{X}}}}+\alpha{\bm{\mathbf{{W}}}}^{\sf T}$
@@
19:         ${\bm{\mathbf{{H}}}}:=\text{Update}\Big{(}{\bm{\mathbf{{G_{W}}}}},{\bm{\mathbf{{Y}}}}_{W}\Big{)}$
@@ 4.2 Practical Considerations for SymNMF via Leverage Score Sampling
This section describes the implementation of Algorithm LvS-SymNMF. We discuss how the leverage scores are computed and the use of a hybrid leverage score sampling scheme which was introduced in [36]. A theoretical analysis of Hybrid Sampling is given later in Section 4.3.2.
@@
Hybrid leverage score sampling samples a subset of rows deterministically and then randomly samples from the remaining rows. We find that Hybrid Sampling is crucial for good performance in our empirical results and we offer a rigorous analysis of its theoretical performance in Section 4.3.2. Hybrid Sampling was proposed and shown to be effective for computing CP decompositions of sparse tensors in [36]. A similar method has been used and theoretically analyzed before for the column subset selection problem [46].
In Hybrid Sampling, a threshold $\tau\in[0,1]$ is used as a hyperparameter. When sampling according to the leverage score distribution, all rows that satisfy $p_{i}\geq\tau$ are deterministically selected. Let the full set of row indices be $\mathscr{I}$, the set that is deterministically included be $\mathscr{I}_{D}$ and $s_{D}=|\mathscr{I}_{D}|$, and the rest be $\mathscr{I}_{R}=\mathscr{I}\setminus\mathscr{I}_{D}$ with $s_{R}=|\mathscr{I}_{R}|$, the remaining indices from which random samples are drawn.
Let ${\bm{\mathbf{{Q}}}}_{A}$ be a $m\times k$ matrix with othonormal columns that is being sampled. Assume without loss of generality that the rows of ${\bm{\mathbf{{Q}}}}_{A}$ permuted conformally to the sets $\mathscr{I}_{D}$ and $\mathscr{I}_{R}$ then the hybrid sampling matrix takes the form
(20) ${\bm{\mathbf{{S}}}}_{H}=\begin{bmatrix}{\bm{\mathbf{{S}}}}_{D}&{\bm{\mathbf{{0}}}}_{s_{D}\times(m-s_{D})}\\ {\bm{\mathbf{{0}}}}_{(s_{R})\times s_{D}}&{\bm{\mathbf{{S}}}}_{R}\end{bmatrix}\in\mathbb{R}^{s_{H}\times m}$
where ${\bm{\mathbf{{S}}}}_{R}\in\mathbb{R}^{s_{R}\times(m-s_{D})}$ and ${\bm{\mathbf{{S}}}}_{D}\in\mathbb{R}^{s_{D}\times s_{D}}$ is a permutation matrix for the deterministically included portion defined as
(21) $({\bm{\mathbf{{S}}}}_{D})_{ji}=\begin{cases}1,\ \text{if row $i$ is the $j$th deterministic sample}\\ 0,\ \text{otherwise}.\end{cases}$
${\bm{\mathbf{{S}}}}_{D}$ is often included in notation to make the deterministic inclusion aspect of the equations explicit. The submatrix ${\bm{\mathbf{{S}}}}_{R}$ is a leverage score sampling matrix as defined in Equation 11 but just of the indices in $\mathscr{I}_{R}$. When sampling for ${\bm{\mathbf{{S}}}}_{R}$, rows that were sampled during the deterministic phase are not considered and the leverage scores are renormalized appropriately. The new leverage scores probabilities are $\tilde{p}_{i}=\frac{l_{i}}{k-\theta}$ where $\theta=\sum_{i\in\mathscr{I}_{D}}l_{i}({\bm{\mathbf{{A}}}})$.
@@ 4.3.1 Leverage Score Sampling for NLS Problems
Error bounds and corresponding sampling complexities for sketching the ordinary least squares (OLS) problem have been previously shown in a number of works [14, 48]. Larsen and Kolda gave a bound and proof structure in their work on computing a randomized CP decomposition [36]. Boutsidis and Drineas [4] considered using the randomized Hadamard transform for solving the NLS problem. Our proof structure generally follows that in [14] for OLS problems.
This section provides a proof of Theorem 2.1 which provides generalized error bounds and sampling complexities for the Nonnegative Least Squares problem (NLS). We note that our results hold for all constrained least squares (CLS) problems as far as the problem remains convex. The NLS problem in Equation 13 is a convex optimization problem but unlike the case of OLS, does not yield a closed form solution. We are concerned only with the case where ${\bm{\mathbf{{A}}}}$ is full rank and overdetermined. The NLS error bound and sample complexity can be derived based on results by Daniels for the perturbation of Convex Quadratic Programs [11] and leverage score sampling for the OLS problem [14, 43, 37, 56].
The proof uses two Structural Conditions (SC’s) such that if both are true then the error bound in Theorem 2.1 for the NLS problem holds. We first discuss the sampling complexity and probability conditions under which these SC’s hold.
Let ${\bm{\mathbf{{A}}}}={\bm{\mathbf{{U}}}}_{A}{\bm{\mathbf{{\Sigma}}}}_{A}{\bm{\mathbf{{V}}}}_{A}^{\sf T}$ be the thin SVD of the coefficient matrix in Equation 13 where ${\bm{\mathbf{{U}}}}_{A}\in\mathbb{R}^{m\times n}$, ${\bm{\mathbf{{\Sigma}}}}_{A}\in\mathbb{R}^{n\times n}$, and ${\bm{\mathbf{{V}}}}_{A}\in\mathbb{R}^{n\times n}$. Let ${\bm{\mathbf{{S}}}}$ be a leverage score sampling matrix for ${\bm{\mathbf{{A}}}}$. The first Structural Condition (SC1) is
(22) $1-\epsilon_{s}\leq\sigma_{i}({\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{S}}}}^{\sf T}{\bm{\mathbf{{S}}}}{\bm{\mathbf{{U}}}}_{A})\leq 1+\epsilon_{s}$
@@
(23) $\|{\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{r}}}}_{nls}-{\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{S}}}}^{\sf T}{\bm{\mathbf{{S}}}}{\bm{\mathbf{{r}}}}_{nls}\|_{2}\leq\sqrt{\epsilon_{r}}\frac{\|{\bm{\mathbf{{r}}}}_{nls}\|_{2}}{\sqrt{2}}$
for some $\epsilon_{r}\in(0,1)$. The leverage score sketching matrix ${\bm{\mathbf{{S}}}}$ satisfies SC1 and SC2 with high probability (given sufficiently many samples). For SC1 this is shown in [56] and for SC2 it can be shown by using Theorem F.2 on the product ${\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{r}}}}_{nls}$ [43]. We include the associated theorems for these statements in Section F.1.
@@
(24) $\min_{{\bm{\mathbf{{x}}}}\in\mathscr{C}}Q({\bm{\mathbf{{x}}}})=\min_{{\bm{\mathbf{{x}}}}\in\mathscr{C}}\frac{1}{2}{\bm{\mathbf{{x}}}}^{\sf T}{\bm{\mathbf{{K}}}}{\bm{\mathbf{{x}}}}-{\bm{\mathbf{{x}}}}^{\sf T}{\bm{\mathbf{{d}}}}$
@@
(25) $\min_{{\bm{\mathbf{{x}}}}\geq 0}\|{\bm{\mathbf{{S}}}}{\bm{\mathbf{{A}}}}{\bm{\mathbf{{x}}}}-{\bm{\mathbf{{S}}}}{\bm{\mathbf{{b}}}}\|_{2}$
@@
(26) $\min_{{\bm{\mathbf{{x}}}}\geq 0}Q({\bm{\mathbf{{x}}}})=\min_{{\bm{\mathbf{{x}}}}\geq 0}\frac{1}{2}{\bm{\mathbf{{x}}}}^{\sf T}{\bm{\mathbf{{A}}}}^{\sf T}{\bm{\mathbf{{A}}}}{\bm{\mathbf{{x}}}}-{\bm{\mathbf{{x}}}}^{\sf T}{\bm{\mathbf{{A}}}}^{\sf T}{\bm{\mathbf{{b}}}},$
(27) $\min_{{\bm{\mathbf{{x}}}}\geq 0}\hat{Q}({\bm{\mathbf{{x}}}})=\min_{{\bm{\mathbf{{x}}}}\geq 0}\frac{1}{2}{\bm{\mathbf{{x}}}}^{\sf T}{\bm{\mathbf{{A}}}}^{\sf T}{\bm{\mathbf{{S}}}}^{\sf T}{\bm{\mathbf{{S}}}}{\bm{\mathbf{{A}}}}{\bm{\mathbf{{x}}}}-{\bm{\mathbf{{x}}}}^{\sf T}{\bm{\mathbf{{A}}}}^{\sf T}{\bm{\mathbf{{S}}}}^{\sf T}{\bm{\mathbf{{b}}}}.$
The objective function give by 27 can be interpreted as a perturbed version of the objective function given by 26 with perturbed parameters $\hat{{\bm{\mathbf{{K}}}}}={\bm{\mathbf{{A}}}}^{\sf T}{\bm{\mathbf{{S}}}}^{\sf T}{\bm{\mathbf{{S}}}}{\bm{\mathbf{{A}}}}$ and $\hat{{\bm{\mathbf{{d}}}}}={\bm{\mathbf{{A}}}}^{\sf T}{\bm{\mathbf{{S}}}}^{\sf T}{\bm{\mathbf{{b}}}}$. We make use of the following, which is a straightforward consequence of the fact that $\nabla f({\bm{\mathbf{{x}}}}^{*})^{\sf T}({\bm{\mathbf{{x}}}}-{\bm{\mathbf{{x}}}}^{*})\geq 0$ for all ${\bm{\mathbf{{x}}}}\in\mathscr{C}$, which is the characterization of a minima (at ${\bm{\mathbf{{x}}}}^{*}$) of a differentiable convex function $f({\bm{\mathbf{{x}}}})$,
Lemma 4.1 ([11, Equation 2.4]).
Convex Quadratic Program Inequality: $Q({\bm{\mathbf{{x}}}})$ and $\hat{Q}({\bm{\mathbf{{x}}}})$ from Equation 26 and Equation 27 satisfy
(28) $(\hat{{\bm{\mathbf{{x}}}}}_{nls}-{\bm{\mathbf{{x}}}}_{nls})^{\sf T}\Big{[}\nabla\hat{Q}(\hat{{\bm{\mathbf{{x}}}}}_{nls})-\nabla\hat{Q}({\bm{\mathbf{{x}}}}_{nls})\Big{]}\leq(\hat{{\bm{\mathbf{{x}}}}}_{nls}-{\bm{\mathbf{{x}}}}_{nls})^{\sf T}\Big{[}\nabla Q({\bm{\mathbf{{x}}}}_{nls})-\nabla\hat{Q}({\bm{\mathbf{{x}}}}_{nls})\Big{]}$
@@
All the tools needed for the proof of Theorem 2.1 have now been established.
Proof 4.2 (Proof of Theorem 2.1).
Substituting the QP formulations of the original and sampled NLS problems into Equation 28, denoting $\hat{{\bm{\mathbf{{x}}}}}_{nls}-{\bm{\mathbf{{x}}}}_{nls}={\bm{\mathbf{{y}}}}$ and parameterizing in terms of the matrix ${\bm{\mathbf{{U}}}}_{A}$ by writing ${\bm{\mathbf{{Z}}}}={\bm{\mathbf{{\Sigma}}}}_{A}{\bm{\mathbf{{V}}}}^{\sf T}_{A}$ and ${\bm{\mathbf{{Z}}}}{\bm{\mathbf{{y}}}}={\bm{\mathbf{{z}}}}$, the right hand side becomes
@@
${\bm{\mathbf{{y}}}}^{\sf T}{\bm{\mathbf{{Z}}}}^{\sf T}{\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{S}}}}^{\sf T}{\bm{\mathbf{{S}}}}{\bm{\mathbf{{U}}}}_{A}{\bm{\mathbf{{Z}}}}{\bm{\mathbf{{y}}}}={\bm{\mathbf{{z}}}}^{\sf T}({\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{S}}}}^{\sf T}{\bm{\mathbf{{S}}}}{\bm{\mathbf{{U}}}}_{A}){\bm{\mathbf{{z}}}}=\|{\bm{\mathbf{{S}}}}{\bm{\mathbf{{U}}}}_{A}{\bm{\mathbf{{z}}}}\|_{2}^{2}\geq\sigma_{min}({\bm{\mathbf{{S}}}}{\bm{\mathbf{{U}}}}_{A})^{2}\|{\bm{\mathbf{{z}}}}\|_{2}^{2}.$
The left hand side of Equation 28 yields